OpenStudy (anonymous):

how to solve the equation of Cos2X+6Sin square X=4

OpenStudy (phi):

\[ \cos(2x) + 6 \sin^2(x) = 4\] ?

OpenStudy (phi):

I would use the double angle formula see http://www.sosmath.com/trig/douangl/douangl.html to put cos(2x) in terms of sin^2(x)

OpenStudy (anonymous):

BUT WHAT IS THE VALUE OF x?

OpenStudy (anonymous):

I CAN FIND COSX IS EQUAL TO + OR - SQUARE ROOT OF COS 2X +2 OVER 6 BUT I CAN'T DETERMINE THE VALUE OF (X)?

OpenStudy (phi):

you need to change cos(2x) into a form with just sin^2(x) can you do that ? see site posted above for the formula

OpenStudy (anonymous):

ok I find sin square (X) IS EQUAL TO 1OVER 2 OR 1\2, THEN X IS EQUAL TO 30 DEGREES?

OpenStudy (phi):

how. can you show your work?

OpenStudy (anonymous):

Ok this the work, I change cos2x by cos square x plus sins square x and I have the complete equation : 1-sin square x + sin square x +6 sin square x =4 , and we have left 6 sins square x=4-1 then sin square =1\2

OpenStudy (anonymous):

by looking for the value of the angle then I get X is equal to 30 degrees of sin square 1\2

OpenStudy (anonymous):

Sorry sin square x = 1\2

OpenStudy (phi):

notice cos2x by cos square x plus sins square x \[ \cos(2x) = \cos^2(x) - \sin^2(x) \] if you use \[ \cos^2(x) = 1 - \sin^2(x) \] you can say \[ \cos(2x) = 1 - 2\sin^2(x)\]

OpenStudy (phi):

it is cos squared MINUS sin squared

OpenStudy (anonymous):

ok thank you I went the other way I will correct it

OpenStudy (anonymous):

now we have x is equal to 48,59 degrees?

OpenStudy (anonymous):

but the sin square doesn't change the value of (x)?

OpenStudy (phi):

can you show your work?

OpenStudy (anonymous):

ok we have as equation :cos2x+6 sin square x =4, I replace cos2x by 1- 2sin square x + 6 sin square x=4. We have left: 4 sin square=4-1 ,then 4 sin square= 3 , we have at the end sin square x= 3\4 and I look for the value of the angle I have 48, 59 degrees in the calculator.?

OpenStudy (phi):

looks ok until the last step \[ \sin^2(x) = \frac{3}{4} \] you want sin(x) not sin squared. take the square root of both sides \[ \sqrt{\sin^2(x) }= \sqrt{\frac{3}{4} } \\ \sin(x) = \frac{\sqrt{3}}{2} \]

OpenStudy (phi):

it should probably be \[ \sin(x) = ±\frac{\sqrt{3}}{2} \]

OpenStudy (anonymous):

Ok I agree with you, I do want sin (x) not sin square (x).That is perfect, you are the best. Sometimes trig is very tricky.

OpenStudy (phi):

you should get x= 60º , 120º, 240º, and 300º

OpenStudy (anonymous):

but I can not get the those values in the calculator, x= root of 3 over 2 and I get error from the machine. I don't know how come this happens?

OpenStudy (phi):

you should do inverse sin (look for the button labeled \(\sin^{-1}\) on sqr(3)/2

OpenStudy (phi):

However, most people memorize certain angles and their sin, cos and tan see http://www.analyzemath.com/trigonometry/special_angles.html

OpenStudy (anonymous):

thank you I do it by hand , I get root of 3 is equal to 1,74 and then I divide it by 2 from I get sinx is equal .86 and I find the angle of the first x which is equal to 59,31 degrees

OpenStudy (phi):

you should do it with the calculator. first do sqr 3 = then divide by 2 you should see 0.86.... then do inverse sin

OpenStudy (anonymous):

yes this what I did

OpenStudy (anonymous):

thank you very much